OpenStudy (anonymous):

If x is an angle in standard position and the terminal side of the angle in on the line y=3x in quadrant III, find: 3cotx-2sinsqx+4tanx+5sqrt10cosx. simplify answer completely... leave as a reduced fraction.

OpenStudy (anonymous):

10 medals to whoever can answer this one.

OpenStudy (anonymous):



OpenStudy (anonymous):

Just find a point on the line in quadrant III, such as (-1, -3) which would give a radial distance out to the point of r=sqrt(10). Then use the definitions\[\sin(x)=v/r; \ \ \cos(x)=u/r;\ \ \tan(x)=v/u\]along with the reciprocal identities.